OpenStudy (mirandaregina):

Find the distance of line AB if A (3,5) and B (-3,-3). A. 8 B. 6 C. 12 D. 10 When I use the distance formula to find the answer, I get 2 but that isn't even an option here. Help?

Nnesha (nnesha):

which formula you used

OpenStudy (amistre64):

3--3 is not 0

OpenStudy (amistre64):

5--3 is not 2

OpenStudy (mirandaregina):

@Nnesha the distance formula

OpenStudy (mirandaregina):

@amistre64 I was following an example that added 3+(-3) and 5+(-3). is that not right?

OpenStudy (amistre64):

the distance between the xs is subtraction, not addition when we want to find the distance between 2 things, a and b, we do: b - a

OpenStudy (amistre64):

d = sqrt( (x1-x2)^2 + (y1- y2)^2 ) ^ ^

OpenStudy (amistre64):

lets 'move' the points so that one of them is at the origin A (3,5) and B (-3,-3). -3,-5 -3,-5 ------------------- A' = 0,0 B' = -9,-8 the distance hasnt changes, just where they are located

OpenStudy (amistre64):

lol, -6,-8

OpenStudy (mirandaregina):

@amistre64 d = (-3 - 3)^2 + (-3 - 5))^2 = (-6)^2 + (-8)^2 = -14 (?)

OpenStudy (amistre64):

sqrt(36 + 64 ) not sqrt(6+8)

Nnesha (nnesha):

how you get negative answer when you take square its mean - 6 times - 6 so its just 36 and same with 8 square

OpenStudy (mirandaregina):

@amistre64 whoops, I forgot to square it lol so d = 10

Nnesha (nnesha):

yep :)

OpenStudy (amistre64):

:)

OpenStudy (mirandaregina):

@amistre64 @Nnesha thanks to the both of you :)

Nnesha (nnesha):

nope mostly thanks to @amistre64